# Few-Shot Adversarial Prompt Learning on Vision-Language Models

### v1

@@ D.2 Case Study: Improving AdvVLP with Our Learning Objective
We further illustrate the adversarial robustness enhancement brought by using our proposed training objective for prompt learning through an intuitive case study. Here, we adapt AdvVLP with both TeCoA loss and our . In Figure 6, our loss brings about an improvement in zero-shot adversarial robustness across ten out of eleven datasets.
@@
Additionally, our training objective results in evident performance gain under few-shot base-to-new generalization, as revealed in Table 7. That is, we not only achieves better base natural accuracy (+11.11%), base PGD-100 accuracy (+6.67%), new natural accuracy (+2.03%), new PGD-100 accuracy (+0.87%), but also maintains superior stability across different trails.
D.3 Ablation for Loss Function Design
In section 3.4, we present our proposed novel training objective tailored for adversarial prompt learning. Our loss is a two-term design, comprising a natural term and an adversarial term. The adversarial term further considers both the consistency and diversity of natural and adversarial features. In practice, we use KL divergence to constrain cross-modal consistency and encourage uni-modal diversity with cosine similarity. In Table 8, we present other possible designs for the loss function and conduct an ablation study under the adversarial base-to-new setting. Our method provide the best robustness across all these loss function settings.
| Natural term | Adversarial term | Base Natural Acc | Base Adv Acc | New Nat Acc | New Adv Acc | |
| Consistency | Diversity | |||||
| ✘ | TeCoA | ✘ | ||||
| ✔ | TeCoA | ✘ | ||||
| ✔ | JS | ✘ | ||||
| ✔ | KL | ✘ | ||||
| ✔ | KL | MSE | ||||
| ✔ | KL | MAE | ||||
| ✔ | KL | Cos | ||||
D.4 Results on Different CLIP Architectures
We provide the adversarial cross-dataset transfer results on another CLIP ViT backbone, ViT-B/16, that is adapted to the proposed method. With the same architectural design, ViT-B/16 divides the input image into smaller patches to better capture and learn image details. This makes ViT-B/16 generally have superior performance over ViT-B/32 in natural image recognition due to its finer granularity, but it also incurs higher computational costs due to longer input sequences. However, when considering tasks involving adversarial robustness, more complex models do not necessarily yield better performance (Wu et al., 2022). We report the results on ViT-B/16 in Table 9. We find that ViT-B/16 does not bring about improved robustness performance, which is due to adversarial prompt learning focusing more on feature alignment and understanding between different modalities rather than detailed features. Therefore, the loss of detailed information resulting from the division of patches in ViT-B/32 is acceptable.
@@ D.4 Results on Different CLIP Architectures
| ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | ||
| Natural Accuracy (%) | 55.40 | 86.90 | 25.00 | 15.00 | 77.40 | 12.50 | 51.90 | 45.80 | 38.50 | 50.00 | 48.70 | 45.84 |
| PGD-100 Accuracy (%) | 24.50 | 63.70 | 13.20 | 10.70 | 45.80 | 4.70 | 16.20 | 22.30 | 10.80 | 20.50 | 19.50 | 23.24 |
D.5 Zero-shot Adversarial Robustness under Different Perturbation Bounds
In this task, we provide adversarial attacks of varying intensities by changing the perturbation bound to test the effectiveness of the model in learning robust representations from different adversarial distributions. Specifically, we set during the training phase respectively, and use the same values during testing as were used in training.
@@ D.5 Zero-shot Adversarial Robustness under Different Perturbation Bounds
D.6 Zero-shot Adversarial Evaluation under Auto-Attack
We consider more powerful Auto-Attack (Croce & Hein, 2020) to evaluate our adapted model. Now that adversarial prompt tuning does not rely on the obfuscated gradient, we use two APGD variants, APGD-CE and APGD-DLR, in Auto-Attack to evaluate our models. In Table 10, we can conclude that Auto-Attack provides a stronger attack and causes varying degrees of performance degradation in each model. Our model still exhibits better resistance to Auto-Attack compared with AdvVP, AdvVLP, and AdvMaPLe. Moreover, by adapting AdvVLP with our learning objective in Appendix D.2, we achieve further performance gain under all three different perturbation bound settings. Note that Auto-Attack uses a fractional attack generator which explores that fraction space by automatically adjusting step size , it serves as a more effective and powerful attacker for zero-shot adversarial robustness evaluation.
| ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@ D.6 Zero-shot Adversarial Evaluation under Auto-Attack
| ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@
| ImageNet | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@
D.7 Detailed Results for Adversarial Few-shot Learning
For adversarial few-shot prompt learning, we plot curves showing how the average natural and robustness accuracy change with varying shot numbers in Figure 2. Here, we present the mean and standard deviation of natural (Table 11)and robustness 12)accuracy for all experimental settings, datasets, and shot numbers, based on our multiple trials. For our proposed method, when given a smaller number of training samples, both the standard deviation of natural accuracy and robustness accuracy are relatively high, indicating that the performance of learning robust representations at this stage depends on the quality of the samples. As the shot number increases, our method exhibits a significant reduction in the standard deviation for both natural and robustness accuracy, demonstrating its ability to stably acquire adversarial robustness.
@@ D.7 Detailed Results for Adversarial Few-shot Learning
D.8 Detailed Results for Adversarial Base-to-New Generalization
For adversarial base-to-new generalization results in 4.2, we further provide the detailed results on each dataset. In Table 13, our method demonstrates preferable learning performance on the majority of datasets. Specifically, in recognition datasets for fine-grained tasks that significantly differ from generic knowledge (DTD, Flowers102, OxfordPets, FGVCAircraft, etc.), our training objective effectively avoids potential failures caused by natural generalization barriers in robustness learning, thus yielding more stable results across multiple trials.
@@ D.8 Detailed Results for Adversarial Base-to-New Generalization
| ImageNet | Base | Natural Acc | ||||||||
@@
D.9 Comparison between Adversarial Text and Vision Prompt
We design most of baseline settings on the top of adversarial vision prompt framework. As a result, most of them belong to cross-modal prompt family, with learnable prompt tokens not only exist in both vision and text input sequences. However, for the sake of completeness, we also consider the design of prompts in a uni-modal context, namely adversarial vision prompts (AdvVP) and adversarial text prompts (AdvTP). In Figure 8, we find that, as the number of available samples increases, both visual and text prompts fail to acquire more robustness-correlated hints for promoting adversarial robustness. However, although it seems difficult for AdvTP to learn proper adversarial text supervision, AdvTP is capable of maintaining preferable natural performance even when only adversarial samples are visible. We believe this can be attributed to the text prompt’s ability to capture semantic information.
Appendix E Reproducibility
@@ Appendix F Limitations
This paper introduces a framework that leverages the architecture of cross-modal prompts to enhance model robustness. This is achieved by adjusting the prompts to learn adversarial-correlated text supervision. However, prompt learning is merely a lightweight strategy for model adaptation, and other parameter-based adaptation methods, such as full-finetuning, are not considered in this work. Furthermore, while our method has empirically shown that a comprehensive consideration of the connections and distinctions between natural and adversarial samples can better learn adversarial text supervision, a systematic theoretical analysis and proof remain elusive. We regard addressing these limitations as our future direction.

### v2

@@ D.4 Case Study: Improving AdvVLP with Our Learning Objective
We further illustrate the adversarial robustness enhancement brought by using our proposed training objective for prompt learning through an intuitive case study. Here, we adapt AdvVLP with both TeCoA loss and our . In Figure 7, our loss improves zero-shot adversarial robustness across ten out of eleven datasets.
@@
Additionally, our training objective results in evident performance gain under few-shot base-to-new generalization, as revealed in Table 10. That is, we not only achieve better base natural accuracy (+11.11%), base PGD-100 accuracy (+6.67%), new natural accuracy (+2.03%), new PGD-100 accuracy (+0.87%), but also maintains superior stability across different trails.
D.5 Results on Different CLIP Architectures
We provide the adversarial cross-dataset transfer results on another CLIP ViT backbone, ViT-B/16, that is adapted to the proposed method. With the same architectural design, ViT-B/16 divides the input image into smaller patches to better capture and learn image details. This makes ViT-B/16 generally have superior performance over ViT-B/32 in natural image recognition due to its finer granularity, but it also incurs higher computational costs due to longer input sequences. However, when considering tasks involving adversarial robustness, more complex models do not necessarily yield better performance [85]. We report the results on ViT-B/16 in Table 11. We find that ViT-B/16 does not bring about improved robustness performance, which is due to adversarial prompt learning focusing more on feature alignment and understanding between different modalities rather than detailed features. Therefore, the loss of detailed information resulting from the division of patches in ViT-B/32 is acceptable.
@@ D.5 Results on Different CLIP Architectures
| ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | ||
| Natural Accuracy | 55.40 | 86.90 | 25.00 | 15.00 | 77.40 | 12.50 | 51.90 | 45.80 | 38.50 | 50.00 | 48.70 | 45.84 |
| PGD-100 Accuracy | 24.50 | 63.70 | 13.20 | 10.70 | 45.80 | 4.70 | 16.20 | 22.30 | 10.80 | 20.50 | 19.50 | 23.24 |
D.6 Zero-shot Adversarial Robustness under Different Perturbation Bounds
In this task, we provide adversarial attacks of varying intensities by changing the perturbation bounds to test the effectiveness of the model in learning robust representations from different adversarial distributions. Specifically, we set during the training phase respectively, and use the same values during testing as were used in training.
@@ D.6 Zero-shot Adversarial Robustness under Different Perturbation Bounds
D.7 Zero-shot Adversarial Evaluation under Auto-Attack
We consider more powerful Auto-Attack [86] to evaluate our adapted model. Now that adversarial prompt tuning does not rely on the obfuscated gradient, we use two APGD variants, APGD-CE and APGD-DLR, in Auto-Attack to evaluate our models. In Table 12, we can conclude that Auto-Attack provides a stronger attack and causes varying degrees of performance degradation in each model. Our model still exhibits better robustness to Auto-Attack compared with AdvVP, AdvVLP, and AdvMaPLe. Moreover, by adapting AdvVLP with our learning objective in Appendix D.4, we achieve further performance gain under all three different perturbation bound settings. Note that Auto-Attack uses a fractional attack generator which explores that fraction space by automatically adjusting step size , it serves as a more effective and powerful attacker for zero-shot adversarial robustness evaluation.
| ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@ D.7 Zero-shot Adversarial Evaluation under Auto-Attack
| ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@
| ImageNet-1K | Caltech101 | DTD | EuroSAT | OxfordPets | FGVCAircraft | Food101 | Flowers102 | StanfordCars | SUN397 | UCF101 | Average | |
@@
D.8 Discussions on Training-time Attack Generation
We adopt Eq. (8) to carry out adversarial attacks during the training process. We did not take into account in Eq. (8). Including in the generation of adversarial samples would make the gradient information focus on the differences between natural and adversarial examples, thereby generating stronger adversarial perturbations with greater differences from the natural examples. However, since we have incorporated this term in our adversarial defense, the model will gradually provide stronger attacks during iterative learning to ensure differences in image features between natural and adversarial samples, making it somewhat redundant in function.
We validate this with the experimental results in Table 13. We can observe that the results of these two methods for generating adversarial attacks are quite similar, indicating that adding in the attack is indeed redundant. Therefore, for the sake of simplicity, we did not include for training-time attack generation.
| Train-time attack generation | Base Nat Acc | Base Adv Acc | New Nat Acc | New Adv Acc |
| 70.52±0.82 | 38.05±2.15 | 49.58±3.55 | 21.86±2.57 | |
| 70.04±0.94 | 38.06±2.23 | 49.56±3.00 | 22.03±2.19 |
D.9 Detailed Results for Adversarial Few-shot Learning
For adversarial few-shot prompt learning, we plot curves showing how the average natural and robust accuracy change with varying shot numbers in Figure 3. Here, we present the mean and standard deviation of natural (Table 14) and robust (Table 15) accuracy for all experimental settings, datasets, and shot numbers, based on our multiple trials. For our proposed method, when given a smaller number of training samples, both the standard deviation of natural accuracy and robust accuracy are relatively high, indicating that the performance of learning robust representations at this stage depends on the quality of the examples. As the shot number increases, our method exhibits a significant reduction in the standard deviation for both natural and robust accuracy, demonstrating its ability to acquire adversarial robustness stability.
@@ D.9 Detailed Results for Adversarial Few-shot Learning
D.10 Detailed Results for Adversarial Base-to-New Generalization
For adversarial base-to-new generalization results in Section 4.2, we further provide the detailed results on each dataset. In Table 16, our method demonstrates preferable learning performance on the majority of datasets. Specifically, in recognition datasets for fine-grained tasks that significantly differ from generic knowledge (DTD, Flowers102, OxfordPets, FGVCAircraft, etc.), our training objective effectively avoids potential failures caused by natural generalization barriers in robustness learning, thus yielding more stable results across multiple trials.
@@ D.10 Detailed Results for Adversarial Base-to-New Generalization
| ImageNet-1K | Base | Natural Acc | ||||||||
@@
D.11 Comparison between Adversarial Text and Vision Prompt
We design most of the baseline settings on the top of the adversarial vision prompt framework. As a result, most of them belong to a cross-modal prompt family, with learnable prompt tokens not only exist in both vision and text input sequences. However, for completeness, we also consider the design of prompts in a uni-modal context, namely adversarial vision prompts (AdvVP) and adversarial text prompts (AdvTP). In Figure 9, we find that, as the number of available examples increases, both vision and text prompts fail to acquire more robustness correlated hints for promoting adversarial robustness. However, although it seems difficult for AdvTP to learn proper adversarial text supervision, AdvTP is capable of maintaining preferable natural performance even when only adversarial examples are visible. We believe this can be attributed to the text prompt’s ability to capture semantic information.
Appendix E Impact Statement
Appendix F Reproducibility
@@ Appendix G Limitations
This paper introduces a framework that leverages the architecture of cross-modal prompts to enhance model robustness. This is achieved by adjusting the prompts to learn adversarial-correlated text supervision. However, prompt learning is merely a parameter-efficient strategy for model adaptation, and other parameter-based adaptation methods, such as full-finetuning, are not considered in this work. Furthermore, while our method has empirically shown that a comprehensive consideration of the connections and distinctions between natural and adversarial examples can better learn adversarial text supervision, a systematic theoretical analysis and proof remain elusive. We regard addressing these limitations as our future direction.
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: We Summarize the main contribution of our paper in the last paragraph of introduction section.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: We discuss the limitations of our paper in Appendix G.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory Assumptions and Proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: Our paper improves adversarial prompt learning from an empirical perspective and does not rely on theoretical proofs and assumptions.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental Result Reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: For datasets, we only use open-source datasets that are publicly available. For codes, we list the original paper of baseline methods in Appendix C.1 with access to their respective code repositories.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental Setting/Details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in the appendix, or as supplemental material.
7.
Experiment Statistical Significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments Compute Resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: We report the compute resources in the first paragraph of Appendix C.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code Of Ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: We have carefully read the NeurIPS Code of Ethics and checked the anonymity of our submission.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader Impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [Yes]
Justification: We discuss the boarder impact of our paper in Appendix E.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: Our paper does not include generative models and typically uses open-source datasets for training and evaluation.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: The creator of assets[28] used in our paper states the license in their repository (MIT License).
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New Assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: Although we will submit the code in the supplementary materials, we will continue to improve the codebase and make it publicly available after the paper is officially accepted. Currently, we have not released any new assets.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and Research with Human Subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: There are no crowdsourcing experiments and research with human subjects under adversarial prompt learning settings.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: There are no crowdsourcing experiments and research with human subjects under adversarial prompt learning settings.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.